# Modal logics of almost sure validities in some classes of euclidean and transitive frames

### v1

@@ 3 Euclidean frames
Let $F=(X,\,R)$ be Euclidean. Then there exists a unique nonempty $U\subseteq X$ such that $R{\upharpoonright}U$ is an equivalence relation and $R\subseteq X\times U.$
@@
To prove the uniqueness, let us assume that $V\subseteq X$ satisfies the consitions. If $a\in V\setminus U$ then $R{}^{-1}(a)=\varnothing,$ so $R$ is not reflexive on $V\setminus U.$ Since $R{\upharpoonright}V$ is an equivalence relation, $V\subseteq U.$ If $a\in U$ then $aRa$, so $(a,\,a)\in X\times V,$ hence $a\in V.$ Then $U=V$, so $U$ is the unique subset of $X$ that satisfies the conditions.
@@
A Euclidean frame $F$ is connected if and only if $F$ is serial and $F{\uparrow}U_{F}$ is a cluster.
Proof 3.4.
Let $F=(X,\,R)$ be connected. Consider $a\not\in U_{F}$. Let $u\in U_{F}$, then $aR^{*}u$ or $uR^{*}a$. But $R[U_{F}]\subseteq U_{F}$, so $R^{*}(u)\subseteq U_{F}$ and $a\not\in R^{*}(u).$ Then $aR^{*}u,$ so $R(a)\neq\varnothing.$ Then $R{\upharpoonright}(X\setminus U_{F})$ is serial; since $R{\upharpoonright}U_{F}$ is reflexive, $R$ is serial. Since $R$ is an equivalence relation on $U_{F},$ $(R^{*}\cup(R^{*}){}^{-1}\cup Id_{X}){\upharpoonright}U_{F}=R{\upharpoonright}U_{F},$ so $R{\upharpoonright}U_{F}=U_{F}\times U_{F},$ thus $F{\uparrow}U_{F}$ is a cluster.
Conversely, suppose that $F=(X,\,R)$ is serial and $R{\upharpoonright}U_{F}$ is a cluster. For any $a\in X,$ $R(a)\subseteq U_{F}$ and $R(a)\neq\varnothing,$ so $aRu$ for some $u\in U.$ Then $aR^{2}v$ for any $v\in U.$ Therefore $R^{2}(X)=U,$ so $R^{2}\cup R^{-2}=X\times X,$ and $F$ is connected.
@@
$\left|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{K5}\right|=2^{\frac{n^{2}}{4}+n+O(\log n)}.$
Proof 3.6.
A connected Euclidean frame $F=(n,\,R)$ is uniquely determined by $U_{F}$ and the family of subsets $R(a)\subseteq U_{F},\,a\in n\setminus U_{F}.$ Then
$\left|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{K5}\right|=\sum_{\varnothing\neq U\subseteq n}2^{|U|(n-|U|)}=\sum_{m=1}^{n}\binom{n}{m}2^{m(n-m)}.$ (11)
@@
$\sum_{m=1}^{n}\binom{n}{m}2^{m(n-m)}$ $\leq\sum_{m=1}^{n}\binom{n}{m}2^{\frac{n^{2}}{4}}=2^{\frac{n^{2}}{4}}\sum_{m=1}^{n}\binom{n}{m}\leq 2^{\frac{n^{2}}{4}}\cdot 2^{n}=2^{\frac{n^{2}}{4}+n};$
$\sum_{m=1}^{n}\binom{n}{m}2^{m(n-m)}$ $\geq\binom{n}{\frac{n}{2}}2^{\frac{n}{2}\left(n-\frac{n}{2}\right)}=\frac{2^{n}}{\sqrt{\pi n/2}}(1+o(1))\cdot 2^{\frac{n^{2}}{4}}=2^{\frac{n^{2}}{4}+n+O(\log n)}.$
@@
For any fixed $r\in\omega,$ $|U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K5})}|>r$ asymptotically almost surely.
Proof 3.8.
By (11), the number of connected Euclidean frames with $|U_{F}|\leq r$ is
$\sum_{U\subseteq n,\,|U|\leq r}2^{|U|(n-|U|)}=\sum_{m=1}^{r}\binom{n}{r}2^{m(n-m)}\leq r\cdot O(n^{r})\cdot 2^{nr}=O(2^{nr}),\,n\to\infty.$
@@
$\operatorname{P}(|U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K5})}|\leq r)=\frac{O(2^{nr})}{\left|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{K5}\right|}=\frac{O(2^{nr})}{2^{\frac{n^{2}}{4}+O(n)}}\to 0,\quad n\to\infty.$
@@
Let $\hat{R}_{n}$ denote the relation of $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K5})$ and $\hat{U}_{n}=U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K5})}.$ Let $r\in\omega$ be fixed. Then a.a.s. there exists a state $a\in n\setminus\hat{U}_{n}$ such that $r<|\hat{R}_{n}(a)|<|\hat{U}_{n}|-r.$
Proof 3.10.
Let $Q_{n,r}$ be the set of connected Euclidean frames $F=(n,\,R)$ such that for any $a\in X\setminus U_{F},$ ${|R(a)|\leq r}$ or $|R(a)|\geq|U_{F}|-r$. Let us estimate $|Q_{n,r}|.$ If we fix $U_{F}\subseteq n$ with $|U_{F}|=m$, then for any $a\not\in U_{F}$ the subset $R(a)\subseteq U_{F}$ can be chosen in $\sum_{k=1}^{r}\left(\binom{m}{k}+\binom{m}{m-k}\right)$ ways. Note that
$\sum_{k=1}^{r}\left(\binom{m}{k}+\binom{m}{m-k}\right)\leq 2r\binom{m}{r}\leq 2r\frac{m^{r}}{r!}\leq 2m^{r}.$
@@
$|Q_{n,r}|\leq\sum_{m=1}^{n}\binom{n}{m}\left(2m^{r}\right)^{n-m}\leq\sum_{m=1}^{n}\binom{n}{m}(2m)^{nr}\\ \leq(2n)^{nr}\sum_{m=1}^{n}\binom{n}{m}\leq(2n)^{nr}2^{n}=O\left(2^{rn\log_{2}n}\right),\quad n\to\infty.$
Finally, we show that $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K5})\not\in Q_{n,r}$ a.a.s. . By Proposition 3.5,
$\operatorname{P}(U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K5})}\in Q_{n,r})=\frac{|Q_{n,r}|}{\left|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{K5}\right|}\leq\frac{O(2^{rn\log_{2}n})}{2^{\frac{n^{2}}{4}+O(n)}}\to 0,\,n\to\infty.$
@@
$\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{K5})=\mathbf{KD5}.$
Proof 3.12.
By Proposition 3.3 every connected Euclidean frame is serial, so $\hat{F}(\operatorname{Con}\operatorname{Fr}\mathbf{K5})\models\mathbf{KD5}$ almost surely.
For the other direction, consider a formula $\varphi\not\in\mathbf{KD5}.$ There exists a finite point-generated Euclidean frame $G=(Y,S)=G{\uparrow}c$ such that $G\not\models\varphi.$ Let $r=|\operatorname{dom}G|.$
Let $F=(X,\,R)$ be a connected Euclidean frame such that
@@
By Propositions 3.7 and 3.9, $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K5})$ has the property (12) a.a.s., so $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K5})\twoheadrightarrow G$, hence $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K5})\not\models\varphi$ a.a.s.
Theorem 3.13.
$\mathbf{K5}^{\mathbf{as}}=\mathbf{KD5}.$
Proof 3.14.
By Theorem 2.7 $\mathbf{K5}^{\mathbf{as}}\subseteq\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{K5})=\mathbf{KD5}.$
It remains to show that $\mathcal{F}_{n}(\mathbf{K5})$ is serial a.a.s. If $F=(n,R)$ is Euclidean and not serial then there exists $U\subsetneq n$ such that $F{\uparrow}U$ is connected and $R{\upharpoonright}(n\setminus U)$ is empty. The subframe $F{\uparrow}U$ may coincide with any frame in $\mathcal{F}_{|U|}\cap\operatorname{Con}\operatorname{Fr}\mathbf{K5}.$ Then, using Proposition 3.5, we get
$\left|\mathcal{F}_{n}\cap\operatorname{Fr}\mathbf{K5}\setminus\operatorname{Fr}\mathbf{KD5}\right|=\sum_{U\subsetneq n}\left|\mathcal{F}_{|U|}\cap\operatorname{Con}\operatorname{Fr}\mathbf{K5}\right|\\ =\sum_{m=0}^{n-1}\binom{n}{m}\left|\mathcal{F}_{m}\cap\operatorname{Con}\operatorname{Fr}\mathbf{K5}\right|\leq n\cdot 2^{n}\cdot\left|\mathcal{F}_{n-1}\cap\operatorname{Con}\operatorname{Fr}\mathbf{K5}\right|\\ =n\cdot 2^{n}\cdot 2^{\frac{(n-1)^{2}}{4}+n+O(\log n)}=2^{n+\frac{n^{2}}{4}-\frac{n}{2}+O(\log n)}=2^{\frac{n^{2}}{4}+\frac{n}{2}+O(\log n)}.$
Then
$\operatorname{P}(\hat{F}_{n}(\mathbf{K5})\not\models\mathbf{KD5})=\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}\mathbf{K5}\setminus\operatorname{Fr}\mathbf{KD5}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}\mathbf{K5}\right|}\leq\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}\mathbf{K5}\setminus\operatorname{Fr}\mathbf{KD5}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{K5}\right|}\\ =\frac{2^{\frac{n^{2}}{4}+\frac{n}{2}+O(\log n)}}{2^{\frac{n^{2}}{4}+n+O(\log n)}}=2^{-\frac{n}{2}+O(\log n)}\to 0,\quad n\to\infty.$
Therefore $\hat{F}_{n}(\mathbf{K5})\models\mathbf{KD5}$ a.a.s., so $\mathbf{KD5}\subseteq\mathbf{K5}^{\mathbf{as}}.$
@@
This theorem follows directly from Proposition 1.3 and Theorem 3.13.
Now we use our method to find $\mathbf{K45}^{\mathbf{as}},\,\mathbf{KD45}^{\mathbf{as}},\,\mathbf{K5B}^{\mathbf{as}},\,\mathbf{S5}^{\mathbf{as}}.$ We follow mostly the same strategy, so we omit some technical steps.
@@
$\mathbf{K45}^{\mathbf{as}}=\mathbf{KD45}^{\mathbf{as}}=\mathbf{KD45}.$
@@
Let us find $\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{K45}).$ Notice that if a frame $F=(X,R)$ is transitive, Euclidean and connected iff $R=X\times U_{F}$. Then such frame is uniquely determined by its cluster $U_{F},$ which must be non-empty if $|X|>1.$ Therefore
$|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{K45}|=\left|\{U\subseteq n:\>U\neq\varnothing\}\right|=2^{n}-1.$
For any fixed $r\in\omega$ we have $r<|U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K45})}|<n$ a.a.s. Indeed, for any $m\leq n$ there are $\binom{n}{m}$ choices for a cluster of size $m$ in $n,$ so
$\operatorname{P}(|U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K45})}|\leq r)=\frac{\sum_{m=1}^{r}\binom{n}{m}}{2^{n}-1}=\frac{O(n^{r})}{2^{n}-1}\to 0,$ $n\to\infty,$
$\operatorname{P}(|U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K45})}|=n)=\frac{1}{2^{n}-1}\to 0,$ $n\to\infty.$
Then for any finite point-generated frame $G=(Y,\,S)\models\mathbf{KD45}$, a.a.s. there exist a p-morphism from some generated subframe of $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K45})$ to $G,$ so  $\operatorname{Log}\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K45})\subseteq\operatorname{Log}G$ a.a.s., thus $\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{K45})\subseteq\mathbf{KD45}.$
By Theorem 2.7 we have $\mathbf{K45}^{\mathbf{as}}\subseteq\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{K45})\subseteq\mathbf{KD45}.$ For the converse inclusion, we show directly that $\hat{F}_{n}(\mathbf{K45})$ is serial a.a.s.
The number of nonserial frames in $\mathcal{F}_{n}\cap\operatorname{Fr}\mathbf{K45}$ is
$\sum_{m=0}^{n-1}\binom{n}{m}|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{K5}|=\sum_{m=0}^{n-1}\binom{n}{m}(2^{m}-1)=O(2^{2n})=o(B_{n}),\,n\to\infty,$
while $|\mathcal{F}_{n}\cap\operatorname{Fr}\mathbf{K45}|\geq B_{n}$ by Proposition 2.1. Then
$\operatorname{P}(\hat{F}_{n}(\mathbf{K45})\text{ is not serial})\leq\frac{o(B_{n})}{B_{n}}\to 0,\,n\to\infty.$
Therefore $\mathbf{KD45}\subseteq\mathbf{K45}^{\mathbf{as}}.$
That $\mathbf{KD45}^{\mathbf{as}}=\mathbf{KD45}$ follows from Proposition 1.3 and Theorem 3.13.
@@
A $\mathbf{K5B}$-frame with set of states $n$ is uniquely determined by a subset $E\subseteq n$ that consists of the irreflexive singletons and an equivalence relation on $n\setminus E.$ Then
$\left|\mathcal{F}_{n}\cap\mathbf{K5B}\right|=\sum_{E\subseteq n}B_{n-|E|}=\sum_{m=0}^{n}\binom{n}{m}B_{n-m}=\sum_{m=0}^{n}\binom{n}{m}B_{n}=B_{n+1}.$

### v2

@@ 3 Euclidean frames
Let $F=(X,\,R)$ be Euclidean. Then there exists a unique subset $U\subseteq X$ (possibly empty) such that $R{\upharpoonright}U$ is an equivalence relation and $R\subseteq X\times U.$
@@
To prove the uniqueness, let us assume that $V\subseteq X$ satisfies the conditions. If $a\in V\setminus U$, then $R{}^{-1}(a)=\varnothing,$ so $R$ is not reflexive on $V\setminus U.$ Since $R{\upharpoonright}V$ is an equivalence relation, $V\subseteq U.$ If $a\in U$, then $aRa$, so $(a,\,a)\in X\times V,$ hence $a\in V.$ Then $U=V$, so $U$ is the unique subset of $X$ that satisfies the conditions.
Definition 3.3.
@@
A serial Euclidean frame $F$ is connected if and only if $F{\uparrow}U_{F}$ is a cluster.
Proof 3.6.
Let $F=(X,\,R)$ be connected. Since $R$ is an equivalence relation on $U_{F},$ $(R^{*}\cup(R^{*}){}^{-1}\cup Id_{X}){\upharpoonright}U_{F}=R{\upharpoonright}U_{F},$ so $R{\upharpoonright}U_{F}=U_{F}\times U_{F},$ thus $F{\uparrow}U_{F}$ is a cluster.
Conversely, suppose that $R{\upharpoonright}U_{F}$ is a cluster. For any $a\in X,$ $R(a)\subseteq U_{F}$ and $R(a)\neq\varnothing,$ so $aRu$ for some $u\in U.$ Then $aR^{2}v$ for any $v\in U.$ Therefore $R^{2}(X)=U,$ so $R^{2}\cup R^{-2}=X\times X,$ and $F$ is connected.
@@
$\left|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{KD5}\right|=2^{\frac{n^{2}}{4}+n+O(\log n)}.$
Proof 3.8.
A frame $F=([n],\,R)\in\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{KD5}$ is uniquely determined by $U_{F}$ and the family of subsets $\{R(a)\mid a\in[n]\setminus U_{F}\}$ where $R(a)\subseteq U_{F}$ and $R(a)\neq\varnothing$ for all $a.$ Then
$\left|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{KD5}\right|=\sum_{\begin{subarray}{c}U\subseteq[n]\\ U\neq\varnothing\end{subarray}}\left(2^{|U|}-1\right)^{n-|U|}=\sum_{m=1}^{n}\binom{n}{m}\left(2^{m}-1\right)^{n-m}.$ (12)
@@
$\sum_{m=1}^{n}\binom{n}{m}(2^{m}-1)^{n-m}\leq\sum_{m=1}^{n}\binom{n}{m}2^{\frac{n^{2}}{4}}=2^{\frac{n^{2}}{4}}\sum_{m=1}^{n}\binom{n}{m}\leq 2^{\frac{n^{2}}{4}}\cdot 2^{n}=2^{\frac{n^{2}}{4}+n};$
$\sum_{m=1}^{n}\binom{n}{m}(2^{m}-1)^{n-m}\geq\binom{n}{\frac{n}{2}}\left(2^{\frac{n}{2}}-1\right)^{n-\frac{n}{2}}\\ =\frac{2^{n}}{\sqrt{\pi n/2}}(1+o(1))\cdot 2^{\frac{n^{2}}{4}}(1+o(1))=2^{\frac{n^{2}}{4}+n+O(\log n)}.$
@@
For any fixed $r\in\omega,$ $|U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})}|>r$ asymptotically almost surely.
Proof 3.10.
By (12), the number of connected $\mathbf{KD5}$-frames with $|U_{F}|\leq r$ is
$\sum_{U\subseteq[n],\,|U|\leq r}2^{|U|(n-|U|)}=\sum_{m=1}^{r}\binom{n}{m}2^{m(n-m)}\leq r\cdot O(n^{r})\cdot 2^{nr}=O(2^{nr}),\,n\to\infty.$
@@
$\operatorname{P}(|U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})}|\leq r)=\frac{O(2^{nr})}{\left|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{KD5}\right|}=\frac{O(2^{nr})}{2^{\frac{n^{2}}{4}+O(n)}}\to 0,\quad n\to\infty.$
@@
Let $\hat{R}_{n}$ denote the relation of $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})$ and $\hat{U}_{n}=U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})}.$ Let $r\in\omega$ be fixed. Then a.a.s. there exists a state $a\in[n]\setminus\hat{U}_{n}$ such that $r<|\hat{R}_{n}(a)|<|\hat{U}_{n}|-r.$
Proof 3.12.
Let $\mathcal{Q}_{n,r}$ be the subset of $F=([n],\,R)\in\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{KD5}$ consisting of all frames $F=([n],R)$ such that for any $a\in[n]\setminus U_{F},$ ${|R(a)|\leq r}$ or $|R(a)|\geq|U_{F}|-r$. Let us estimate $|\mathcal{Q}_{n,r}|.$ If we fix $U_{F}\subseteq[n]$ with $|U_{F}|=m$, then for any $a\not\in U_{F}$ the nonempty subset $R(a)\subseteq U_{F}$ can be chosen in $\sum_{k=1}^{r}\left(\binom{m}{k}+\binom{m}{m-k}-1\right)$ ways. Note that
$\sum_{k=1}^{r}\left(\binom{m}{k}+\binom{m}{m-k}-1\right)\leq 2r\binom{m}{r}\leq 2r\frac{m^{r}}{r!}\leq 2m^{r}.$
@@
$|\mathcal{Q}_{n,r}|\leq\sum_{m=1}^{n}\binom{n}{m}\left(2m^{r}\right)^{n-m}\leq\sum_{m=1}^{n}\binom{n}{m}(2m)^{nr}\\ \leq(2n)^{nr}\sum_{m=1}^{n}\binom{n}{m}\leq(2n)^{nr}2^{n}=O\left(2^{rn\log_{2}n}\right),\quad n\to\infty.$
Finally, we show that $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})\not\in\mathcal{Q}_{n,r}$ a.a.s. By Proposition 3.7,
$\operatorname{P}(U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})}\in\mathcal{Q}_{n,r})=\frac{|\mathcal{Q}_{n,r}|}{\left|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{KD5}\right|}\leq\frac{O(2^{rn\log_{2}n})}{2^{\frac{n^{2}}{4}+O(n)}}\to 0,\,n\to\infty.$
@@
$\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})=\mathbf{KD5}.$
Proof 3.14.
By Proposition 3.5 every connected Euclidean frame is serial, so $\hat{F}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})\models\mathbf{KD5}$ almost surely.
For the other direction, consider a formula $\varphi\not\in\mathbf{KD5}.$ There exists a finite point-generated $\mathbf{KD5}$-frame $G=(Y,S)=G{\uparrow}c$ such that $G\not\models\varphi.$ Let $r=|\operatorname{dom}G|.$ Let $F=(X,\,R)$ be a connected $\mathbf{KD5}$-frame such that
@@
By Propositions 3.9 and 3.11, $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})$ has the property (13) a.a.s., so $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})\twoheadrightarrow G$, hence $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})\not\models\varphi$ a.a.s.
@@
By Theorem 1.1, $\mathbf{KD5}\subseteq\mathbf{KD5}^{\mathbf{as}}$. For the other direction, observe that the cluster $([n],\,[n]\times[n])$ is a connected $\mathbf{KD5}$-frame for all $n\in\omega$, thus $\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{KD5}\neq\varnothing$ for all $n\in\omega.$ By Theorem 2.9 $\mathbf{KD5}^{\mathbf{as}}\subseteq\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})=\mathbf{KD5}.$
Now we use our method to find $\mathbf{KD45}^{\mathbf{as}},\,\mathbf{K5B}^{\mathbf{as}},\,\mathbf{S5}^{\mathbf{as}}.$ We follow mostly the same strategy, so we omit some technical steps.
@@
$\mathbf{KD45}^{\mathbf{as}}=\mathbf{KD45}.$
@@
Let us find $\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45}).$ Notice that a Euclidean frame $F=(X,R)$ is serial, transitive and connected iff $U_{F}\neq\varnothing$ and $R=X\times U_{F}$. Then such frame is uniquely determined by its cluster $U_{F}\neq\varnothing$. Therefore
$|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{KD45}|=\left|\{U\subseteq[n]:\>U\neq\varnothing\}\right|=2^{n}-1.$
For any fixed $r\in\omega$ we have $r<|U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})}|<n$ a.a.s. Indeed, for any $m\leq n$ there are $\binom{n}{m}$ choices for a cluster of size $m$ in $n,$ so
$\operatorname{P}(|U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})}|\leq r)=\frac{\sum_{m=1}^{r}\binom{n}{m}}{2^{n}-1}=\frac{O(n^{r})}{2^{n}-1}\to 0,$ $n\to\infty,$
$\operatorname{P}(|U_{\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})}|=n)=\frac{1}{2^{n}-1}\to 0,$ $n\to\infty.$
Then for any finite point-generated frame $G=(Y,\,S)\models\mathbf{KD45}$, a.a.s. there exist a p-morphism from some generated subframe of $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})$ to $G,$ so  $\operatorname{Log}\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})\subseteq\operatorname{Log}G$ a.a.s., thus $\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})\subseteq\mathbf{KD45}.$
By Theorem 2.9 we have $\mathbf{KD45}^{\mathbf{as}}\subseteq\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})\subseteq\mathbf{KD45}.$ The converse inclusion follows from Theorem 1.1.
@@
A $\mathbf{K5B}$-frame with set of states $n$ is uniquely determined by a subset $E\subseteq[n]$ that consists of the irreflexive singletons and an equivalence relation on $n\setminus E.$ Then
$\left|\mathcal{F}_{n}\cap\mathbf{K5B}\right|=\sum_{E\subseteq[n]}B_{n-|E|}=\sum_{m=0}^{n}\binom{n}{m}B_{n-m}=\sum_{m=0}^{n}\binom{n}{m}B_{n}=B_{n+1}.$